The length of a table plus 1/5 of its length
is 36 inches. How long is the table, in inches?
An outside wall is 5 7/8 inches thick. The wall consists of 1/2 inch of drywall and 3 3/4 inches of Insalation, and 5/8 inch of wall sheathing. The remaining thickness is siding. How thick is the siding?

Answer by checkley79(3341)

X+X/5=36
(5X+X)/5=36
6X/5=36
6X=5*36
6X=180
X=180/6
X=30 ANS.
PROOF:
30+30/5=36
30+6=36
36=36
---------------------------------
5 7/8 = 1/2 + 3 3/4 + 5/8 + X
47/8=1/2+15/4+5/8+X LCD=8
47/8=(4+2*15+5)/8+X
47/8=(4+30+5)/8+X
47/8=39/8+X
47/8-39/8=X
8/8=X
1=X ANS. 1 INCH OF SIDING.